OpenStudy (anonymous):

lim as x approaches 0 for f(x)=ln(x/(x-1)

OpenStudy (turingtest):

\[\frac{\ln x}{x-1}\]??

OpenStudy (anonymous):

not quite, natural log of the entire fraction

OpenStudy (turingtest):

\[f(x)=\ln(\frac x{x-1})\]yes?

OpenStudy (anonymous):

yep

OpenStudy (nenadmatematika):

-ininity

OpenStudy (anonymous):

how did you get that?

OpenStudy (turingtest):

long division\[\ln(1+\frac1{x-1})\]x goes to zero and we get\[\ln(1+(-1))=\ln 0=-\infty\]

OpenStudy (nenadmatematika):

ln is continous, so lim and ln can change places, so when you do that, and than plug in x=0 you get ln0 which is the result turing test wrote

OpenStudy (anonymous):

ok that makes sense. Thank you

OpenStudy (nenadmatematika):

you're welcome :D